# LLMs in Biomedicine: A study on clinical Named Entity Recognition

### v1

@@ 7 Conclusion
In this paper, we explored LLMs by customizing various prompting techniques specifically for the biomedical NER. Through a detailed comparative analysis, we have highlighted the vital role of ICL and the selection of contextually pertinent examples, demonstrating the benefits to clinical NER tasks. Moreover, our investigation into incorporating external clinical knowledge resulted in a novel data augmentation approach, considerably advancing the capabilities of zero-shot clinical NER with LLMs.
@@
While we have shown the potential of enhancing LLM performance for clinical NER, the experiments in this paper are limited in two aspects mainly due to computational constraints. (1) TANL uses a straightforward text-to-text format while DICE uses additional descriptions. Future work could attempt to simplify DICE or combine it with TANL. Ablation studies on components of each format could help researchers design new prompt formatting strategies. (2) Our RAG-based method exclusively utilizes UMLS as the knowledge base which is limited in its vocabulary. For medical terms not covered by UMLS, we did not augment the input text, though other knowledge bases such as Wikipedia could potentially serve as an alternative.
8 Aknowledgment
This work was supported in part by Optum. We would like to thank the reviewers for their feedback to improve this research work.
@@ Appendix A Appendix
Fine-tuning setting for Llama
We fine-tuned Llama2-7B for three epochs and maintained a batch size of 16, learning rate of 2e-4, and cap the maximum sequence length at 512, truncating any sequences that exceeded this limit. The LoRA dropout rate is adjusted to 0.1, and the LoRA $\alpha$ and rank parameters are also set at 16 and 32 respectively. The training was done on 4 NVIDIA Tesla V100 GPUs for an approximate of 2 days.
TANL and DICE more examples
In this section, we visualize two examples of each format for NCBI-disease and BC2GM datasets for demonstration.

### v2

@@ 7 Conclusion
We explored LLMs for biomedical NER by customizing various prompting techniques. Through a detailed comparative analysis, we highlighted the vital role of ICL and the selection of contextually pertinent examples with biomedical text encoders for biomedical NER tasks. Moreover, our investigation into incorporating external medical knowledge resulted in a novel data augmentation approach, considerably advancing the capabilities of zero-shot biomedical NER with LLMs.
@@
While we have shown the potential of enhancing LLM performance for biomedical NER, the experiments in this paper are limited in two aspects mainly due to computational constraints. (1) TANL uses a straightforward text-to-text format while DICE uses additional descriptions. Future work could attempt to simplify DICE or combine it with TANL. Ablation studies on components of each format could help researchers design new prompt formatting strategies. (2) Our RAG-based method exclusively utilizes UMLS as the knowledge base, though it is limited in its vocabulary. For medical terms not covered by UMLS, we did not augment the input text. Other knowledge bases such as Wikipedia could serve as an alternative.
@@ Appendix A Appendix
A.1 TANL/DICE more examples
In Fig 4-8, we visualize two examples of each format for NCBI-disease and BC2GM datasets for more demonstration.
A.2 Benchmark datasets
We studied LLMs on I2b2, NCBI-disease, and BC2GM dataset. In the following, we provide some details about each.
I2B2: I2B2 is a collection of annotated clinical records that are used primarily for Clinical NER.The task involves identifying clinical terms such as medical problems, treatments, and tests from patient records. The dataset typically includes a large number of annotated clinical narratives that are de-identified to protect patient confidentiality. This makes it a rich resource for training and testing NER models.
NCBI-disease: This dataset is specifically curated for disease name recognition and normalization in biomedical texts. It comprises abstracts from PubMed annotated for disease mentions and linked to the NCBI disease database. The corpus is relatively smaller compared to i2b2 but is densely annotated, providing high-quality, fine-grained annotations of disease entities, which are crucial for models aimed at medical literature.
BC2GM: This dataset focuses on the recognition of gene and gene product mentions in PubMed abstracts that is a suitable dataset for biological NER. The BC2GM dataset is extensively annotated to include a wide range of gene and gene product mentions, reflecting the complex and varied ways these entities are referred to in scientific literature.
A.3 PEFT setting of Llama for fine-tuning
We fine-tuned Llama2-7B on the entire training set of each dataset for three epochs and maintained a batch size of 16, learning rate of 2e-4, and cap the maximum sequence length at 512, truncating any sequences that exceeded this limit. The LoRA dropout rate is adjusted to 0.1, and the LoRA $\alpha$ and rank parameters are also set at 16 and 32 respectively. The training was done on 4 NVIDIA Tesla V100 GPUs for approximately 24, 12, and 63 hours for I2B2, NCBI-disease, and BC2GM respectively.
A.4 Few-shot and Zero-shot performances with Confidence Interval
We introduced both few-shot and zero-shot settings to comprehensively evaluate the versatility and generalization capabilities of our study across different levels of data availability. While it’s true that the performance in the zero-shot setting is generally lower compared to the few-shot setting, this approach offers valuable insights into the model’s behavior when no training examples are provided. The zero-shot setting, leveraging techniques like Retrieval-Augmented Generation (RAG), demonstrates the model’s potential to utilize pre-existing knowledge embedded in its parameters and external sources effectively. This is particularly important for scenarios where labeled data is scarce or unavailable, making zero-shot learning a critical area of study to ensure broader applicability of the model in real-world applications. Moreover, the inclusion of both methodologies allows us to highlight the performance trade-offs and strengths of the model under different instructional paradigms, contributing to a more robust and nuanced understanding of its capabilities. We ran all experiments with different random seeds and reported the full results of Table 2, 3, and 4 with confidence Intervals In Tables 6, 7, and 8.
A.5 UMLS detail
In Fig 8, we visualize the process by which potential words suggested by the LLM are searched within the UMLS and demonstrate how the input is augmented to enhance zero-shot prompting in LLMs.
@@ A.5 UMLS detail
Mention/Token
Model input-output format I2B2 NCBI-disease BC2GM
GPT-3.5-turbo DICE 41.2 $\pm$ 0.2 /50.0 $\pm$ 0.1 45.3 $\pm$ 0.2 /62.0 $\pm$ 0.3 43.3 $\pm$ 0.5 /55.6 $\pm$ 0.4
TANL 52.9 $\pm$ 0.3/59.7 $\pm$ 0.4 46.5 $\pm$ 0.5/51.3 $\pm$ 0.5 39.1 $\pm$ 0.4/50.8 $\pm$ 0.5
GPT-4 DICE 58.8 $\pm$ 0.4 /70.1 $\pm$ 0.3 68.1 $\pm$ 0.9/77.8 $\pm$ 1.1 57.1 $\pm$ 0.6/67.9 $\pm$ 0.5
TANL 61.9 $\pm$ 0.3/73.5 $\pm$ 0.5 67.5 $\pm$ 0.8/70.0 $\pm$ 0.6 56.4 $\pm$ 0.2/69.6 $\pm$ 0.3
Table 5: TANL vs. DICE format with GPT-3.5-turbo/GPT-4 with confidence intervals
Model KATE vs RS I2B2 NCBI-disease BC2GM
M/T M/T M/T
GPT-3.5-turbo (ICL) RS 52.9 $\pm$ 0.3 / 59.7 $\pm$ 0.4 46.6 $\pm$ 0.5 / 51.3 $\pm$ 0.5 39.1 $\pm$ 0.4 / 50.8 $\pm$ 0.5
[…]
BioClinicalBERT 67.0 $\pm$ 0.6/78.9 $\pm$ 0.5 67.6 $\pm$ 0.1/78.8$\pm$ 0.1 60.9 $\pm$ 0.7 /72.0 $\pm$ 0.5
[…]
SimCSE 65.2 $\pm$ 0.2 / 76.1 $\pm$ 0.3 61.6 $\pm$ 0.4 / 76.1 $\pm$ 0.3 57.8 $\pm$ 0.5 / 68.8 $\pm$ 0.4
Hu et al. (2024b) 49.3/ - - -
GPT4 (ICL) RS 67.7 $\pm$ 0.3 / 73.5 $\pm$ 0.5 62.6 $\pm$ 0.8 /70.0 $\pm$ 0.6 59.2 $\pm$ 0.2 / 69.6 $\pm$ 0.3
[…]
BioClinicalBERT 81.7 $\pm$ 0.4/88.1 $\pm$ 0.3 79.3 $\pm$ 0.4/88.0 $\pm$ 0.3 71.9 $\pm$ 0.3 / 79.4 $\pm$ 0.3
[…]
SimCSE 79.6 $\pm$ 0.5 / 86.6 $\pm$ 0.4 77.3 $\pm$ 0.5 / 86.5 $\pm$ 0.8 69.9 $\pm$ 0.8 / 77.9 $\pm$ 0.5
Hu et al. (2024b) 59.3/ - - -
BioBERT fully supervised - /87.3 - /89.1 - /83.8
BioClinicBERT fully supervised - /87.7 - /89.0 - /81.7
BioClinicRoBERTa fully supervised - /89.7 - /89.0 - /87.0
Table 6: Random example selection (RS) vs. KATE with medical/non-medical encoders vs. fully supervised models with Mention/Token-level (M/T) analysis. KATE significantly outperforms random sampling in all settings, and LMs pre-trained on the biomedical text outperform strong, general domain encoders. HunFlair is added to the paper
Model I2B2 NCBI-disease BC2GM
M/T M/T M/T
Performance GPT-3.5-turbo w/ KATE 67.0 $\pm$ 0.6 / 78.9 $\pm$ 0.5 68.0 $\pm$ 0.3 / 78.8 $\pm$ 0.1 61.6 $\pm$ 0.1 / 72.5 $\pm$ 0.6
GPT4 w/ KATE 81.7 $\pm$ 0.4 / 88.4 $\pm$ 0.6 79.3 $\pm$ 0.4 / 88.3 $\pm$ 0.8 72.4 $\pm$ 0.6 / 80.7 $\pm$ 0.4
Llama2-7B 61.2 $\pm$ 1.8 / 76.2 $\pm$ 1.3 80.4 $\pm$ 0.9/ 91.3 $\pm$ 1.1 68.1 $\pm$ 1.4 / 75.1 $\pm$ 1.3
Cost (T+I) GPT3.5-turbo w/ KATE (\$0.35) (\$0.11) (\$1.34)
GPT4 w/ KATE (\$10.42) (\$3.12) (\$40.13)
Llama2-7B (\$47.85+\$7.4) (\$23.5+\$1.2) (\$69.7+\$12.9)
Model I2B2 NCBI-disease BC2GM
M/T M/T M/T
UniversalNER Zhou et al. (2023) 40.4/ - 60.4/ - 47.2/ -
Rohanian et al. (2023) w/ GPT-3.5 - 33.4 / - 32.0 / -
Hu et al. (2024b) w/ GPT-3.5-turbo 39.3/ - - -
Hu et al. (2024b) w/ GPT-4 52.6/ - - -
HunFlair Weber et al. (2021) 0.0 / 0.0 24.8 / 36.1 28.2 / 22.7
[…]
GPT-3.5-turbo w/ DiRAG 43.0 $\pm$ 0.9 / 55.7 $\pm$ 1.5 44.7 $\pm$ 0.5 / 50.0 $\pm$ 2.1 30.45 $\pm$ 1.6 / 22.5 $\pm$ 2.1
[…]
GPT-4 w/ DiRAG 53.1 $\pm$ 1.1 / 62.8 $\pm$ 1.2 61.0 $\pm$ 0.6 /66.2 $\pm$ 0.5 51.1 $\pm$ 2.0 / 55.0 $\pm$ 2.2
Table 8: Full results of Zero-shot NER with GPT-3.5-turbo and GPT-4, w/ and w/o DiRAG, and their comparision with SOTA. Our method improved zero-shot NER significantly for I2B2 and NCBI-disease datasets.
[…]
Figure 8: UMLS search. The GPT model is prompted for a simpler task of identifying all words that could potentially be a named entity. Then, the retrieved information from UMLS will augment the original input text for recalling the LLM